# SelfAI: Building a Self-Training AI System with LLM Agents

### v1

@@ 1 Main
To address these issues, we propose SelfAI, a unified multi-agent self-training pipeline for continuous adaptability and transparent collaboration. As shown in Fig. 1, SelfAI forms a closed loop that integrates user intent, cognitive reasoning, and experimental orchestration. Upstream, a User Agent converts high-level objectives and exploratory questions into standardized experiment configurations. These configurations enable the Cognitive Agent to analyze performance metrics, conduct reasoning over historical outcomes, and adjust the search trajectory, driving gradual improvement [54, 55]. A resource-aware Experiment Manager ensures efficient execution by handling resource scheduling, environment provisioning, adaptive checkpointing, and comprehensive experiment logging. During high-level parameter optimization, the Cognitive Agent and Experiment Manager jointly estimate the potential benefit of continued training and apply an optimal-stopping criterion to terminate unpromising trials. To evaluate scientific reasoning quality, we introduce two complementary evaluation metrics, Score (Discovery Efficiency) and $\text{AUP}_{D}$ (Area Under the Performance Diversity). Score aggregates, across tasks, the normalized improvement over the search space together with penalties for discovering good configurations late and for stopping far from the best-found point. $\text{AUP}_{D}$ explicitly encodes how broadly a solver explores the search space by summarizing the performance-diversity tradeoffs across the entire trajectory, enabling detailed analysis of exploration behavior and stopping decisions in long-term searches. Together, these metrics enable quantitative assessment of exploration behavior, reasoning structure, and stopping decisions in long-term autonomous experiments.
In summary, SelfAI provides a continuously self-training system that autonomously iterates through cycles of intent-centric understanding, trajectory-aware reasoning, and adaptive strategic planning, enabling sustained performance gains with minimal human intervention. Through checkpoint tolerance, zero-code parallelization, and dynamic resource management, SelfAI substantially reduces redundant computation, improves resource efficiency, and accelerates search convergence in long-horizon scientific workflows. In addition, SelfAI integrates seamlessly with modern MLOps pipelines [56, 57, 58], facilitating collaborative experiment coordination, end-to-end traceability of model evolution, and robust privacy protection for sensitive configurations and user data. Collectively, these capabilities make SelfAI a practical, scalable, and operationally mature platform for rapid scientific discovery. The key contributions of this paper are given as follows:
  1. 1.
    Autonomous Scientific Discovery System: We design a cohesive framework of specialized LLM agents, including the User Agent, Cognitive Agent Framework, and Experiment Manager, which collaboratively translate high-level research objectives into iteratively refined experimental workflows, including trajectory analysis, hypothesis generation, and best stopping decision-making.
  2. 2.
    Novel Evaluation Metrics and Optimal Stopping: We propose two new metrics, Score (discovery efficiency) and $\text{AUP}_{D}$ (area under the performance–diversity curve), to jointly quantify the efficiency and diversity of hyperparameter exploration. An embedded optimal stopping criterion automatically terminates unpromising trials, effectively balancing exploration and exploitation while minimizing computational overhead.
  3. 3.
    Comprehensive Validation of LLM-based Reasoning in Scientific Discovery: We evaluate SelfAI across diverse tasks, including regression, sentiment analysis, computer vision, medical imaging, and drug discovery. The results show that SelfAI effectively improves upon excellent baseline models in both results exploration and performance.
2 Performance of SelfAI on Benchmark Data
Figure 2: Scores among all solvers across different tasks to measure the best stopping criterion.
Figure 3: Diverse Metrics ($\text{AUP}_{D}$) among all solvers across different tasks to evaluate trajectory diversity.
Scientific discovery is driven by diverse physical environments and user intents. To effectively reason across diverse scientific discovery tasks, our approach integrates multiple LLM agents and a suite of tools: User Agent, Cognitive Framework, and Experiment Management, addressing the complex reasoning and planning problem of scientific discovery. In our experiments, we evaluate the SelfAI across 6 domains and 12 tasks. The compared methods contain different suites of LLMs, such as OpenAI-o3 [1], Llama3.3 [59], Qwen2.5 [2], and DeepSeek-R1 [3]. We used grid search and the Tree Parzen Estimator (TPE) optimizer [60, 61] (referred to as BS), which applies a modified Bayesian inference method using a spiral search. All agents are tested on four Nvidia A100 GPUs. A public implementation of SelfAI will be available on the project website. We also provide reasoning processes for reproducibility.
[…]
Benchmarks. We collect benchmarks composed of multiple tasks of 6 primary categories, designed to simulate challenging scenarios in scientific discovery. The benchmark requires the solver to engage in scientific reasoning by analyzing experiments, interpreting observations, and adapting strategies, uncovering novel patterns, functional relationships, or optimal experimental configurations. Compared with hyperparameter search, these processes reflect the inductive reasoning essential to real-world scientific inquiry. Therefore, this benchmark assesses the model’s reasoning ability, not just its numerical optimization capabilities. All benchmark data is collected through actual experimental runs, and some results are reported in published literature, spanning both discrete and continuous hyperparameter search spaces, from low-dimensional to high-dimensional settings (see Appendix B for more details).
Existing methods [24, 45, 49] explore final performance or medal rates using reward functions and prompts specifically designed for innovative idea generation, code generation, and pass@K strategies. In the line of LLM-driven optimization, LLM4EO [50] uses LLMs exclusively to synthesize and evolve evolutionary operators that determine which parts of a candidate solution should be modified during the search. The LLM is used to generate and update operators to shape the local search behavior of the evolutionary algorithm. These implementations do not focus on scientific reasoning across trials during the exploration process, making direct comparisons with our benchmark infeasible. Consequently, we used the readily accessible GPT4-o3-mini model and compared SelfAI with LLM and LLM-ES methods on 11 small/medium-sized tasks across different domains in Fig. 1c. The LLM and LLM-ES methods are fine-tuned by using the structured prompts in Supplementary Section D. The LLM solver uses a search prompt that recommends new trials based on prior results, while LLM-ES augments this with an early-stopping module that decides whether to terminate exploration. Fig. 1c presents that our SelfAI outperforms LLM and LLM-ES in all four metrics, validating SelfAI’s effectiveness in integrating domain knowledge to accelerate scientific discovery. Furthermore, we focus on the reasoning capability of the original LLM model because it is difficult to be fair when comparing different methods. Fig. 1d shows that SelfAI promptly focuses on exploiting higher-performance regions while maintaining a balanced distribution for effective exploration. Figs. 2 and 3 summarize the benchmark performance based on these two metrics. The benchmarks show that SelfAI consistently achieves excellent results across various domains and tasks for different LLM models. Detailed comparisons covering Score, $\text{AUP}_{D}$, Best-Time, Stop-Time, and Hit Rate are provided in Supplementary Table 3.
Failure cases. Several key failure modes of LLM-based solvers are shown in Supplementary Section B.1. First, the hit rate for identifying the best result varies substantially across models, with some solvers (particularly DeepSeek-R1 series) failing to reach the global optimum due to premature stopping. Second, limited context windows cause incomplete reasoning over long trajectories, preventing solvers from fully incorporating earlier experimental signals. Third, computational fragility leads to non-monotonic performance under seemingly minor perturbations in reasoning strategy. These findings indicate that while SelfAI improves overall efficiency, reasoning-driven exploration remains sensitive to solver stability and stopping behavior.
2.1 Machine Learning
Boston house pricing prediction. We evaluate SelfAI on the Boston housing price prediction task [62] using a random forest regression model. A total of 162 trials were conducted across a search space defined by five hyperparameters: i.e., $\text{n-estimators}=[100,200,300]$, $\text{max-depth}=[\text{None},10,20]$, $\text{min-samples-split}=[2,5,10]$, min-samples-leaf=[1,2,5] and $\text{max-features}=[\text{``sqrt'',``log2''}]$). As shown in Figs. 2 and 3 (and Supplementary Table 4), GPT-4o-mini achieves by far the highest Score (0.9811) and the 1st ranking, indicating outstanding optimization efficiency. Within the DeepSeek-R1 family, the 32B and 7B variants secure 2nd and 3rd places respectively, demonstrating the best balance between rapid convergence (low Best-Time) and reasonable search diversity among open-source models. In contrast, DeepSeek-R1-70B and Llama3.3-70B exhibited consistent exploration but lacked effective stopping criteria, leading to longer convergence times. Notably, nearly all models achieved an identical and high Best Result ( 0.841), indicating that while most can find a correct solution, they differ substantially in optimization efficiency and reliability.
Sentiment analysis. We perform experiments for sentiment analysis that focus on identifying opinions, emotions, and attitudes expressed in text. All experiments are found in [63] with standard experimental settings, where the pre-trained Word2Vec embeddings [64] are used as input features. An LSTM network is then employed to model sentence sequences, converting them into a multi-class classification problem. This setup serves to evaluate the agent’s reasoning and optimization capabilities in textual domains. As shown in Supplementary Table 5, GPT-4o-mini outperforms all other LLMs, achieving the highest Score (0.8824) by discovering the optimal configuration extremely early while maintaining reasonable search diversity. DeepSeek-R1 series and Qwen2.5 series exhibit significantly lower Scores, with many failing to converge quickly or occasionally missing the global optimum. The largest frontier model GPT-4o also underperforms markedly (Score 0.2745, rank 11), suggesting that scale alone is insufficient for effective iterative hyperparameter reasoning in recurrent neural architectures.
2.2 Scientific Computing for Image Completion
For scientific computing fields, we selected the tensor decomposition method [65]. The method is an important tool for high-dimensional data analysis and is crucial in applications such as data compression [66], computational acceleration [67], and multi-modal data fusion [28]. All LLMs are provided with identical mathematical knowledge about TW decomposition generated by GPT-4o. As shown in Supplementary Table 6, Qwen2.5 (7b and 14b) and GPT4-o3-mini perform well and rank highly. Qwen2.5-72b, on the other hand, may overly rely on general reasoning capabilities and fail to balance mathematical knowledge and reasoning capabilities. In the DeepSeek series, although the DeepSeek-R1-7b model can’t find the optimal solution, its search strategy was acceptable with a moderate Score, demonstrating its ability to perform mathematical reasoning. DeepSeek-R1-70b and GPT-4o, despite excellent search diversity and early stopping, received a mediocre score, reflecting that their exploration strategy was not well aligned with the optimization goal of tensor decomposition, possibly failing to find the optimal performance between data compression and computational efficiency.
2.3 Computer Vision
A. SIREN We employed SIREN (Sinusoidal Representation Networks) [68] to evaluate our framework on image segmentation and denoising tasks. Leveraging sine activations and coordinate-based inputs, SIREN excels at representing high-frequency signals through continuous implicit representations, making it widely used in scientific computing and physics-based problems. However, its performance is highly sensitive to hyperparameters (e.g., learning rate and regularization strength, etc.), requiring careful tuning per dataset. The unsupervised nature of SIREN further increases the risk of training instability or divergence with improper settings, making it a strong test case for SelfAI’s hyperparameter optimization capability.
Fig. 4 illustrates hyperparameter search trajectories for image segmentation using SIREN (see also Supplementary Figs. 8, 9 and 10), where the surface, smoothed from the original steep, multi-peak data, reveals distinct search behaviors. Given the same three initial points, the tree-structured Bayesian optimizer (BS) follows a spiral trajectory that underutilizes promising starting regions and fails to explore broadly, often converging to local minima. In contrast, LLM-based optimizers infer trends from evaluated points, incorporate causal reasoning, and explore more broadly to efficiently locate the global optimum. They also monitor progress and halt when improvements plateau, a capability absent in traditional methods.
Quantitative results (Supplementary Tables 78) reveal distinct task-dependent performance profiles. In segmentation, DeepSeek-R1-7B achieves the highest Score (0.693) and ranks first, followed by Qwen2.5-7B and GPT-4-o3-mini. Larger models such as Qwen2.5-72B and Llama-3.3-70B rank only mid-range, indicating that scale alone does not ensure effective optimization. In denoising, the landscape shifts: Qwen2.5-72B delivers the best performance (Score 0.761), while DeepSeek-R1-14B and DeepSeek-R1-32B jointly occupy second place. Notably, models that perform well in segmentation tasks do not necessarily perform well in denoising tasks, suggesting that the effectiveness of the solver largely depends on the degree of matching between its inference strategy and the task.
B. Image Classification We also benchmark two commonly used supervised learning methods in computer vision: Masked Autoencoder (MAE) [69] and ResNet [70]. These represent fundamentally different learning paradigms where hyperparameter optimization plays a crucial role in achieving the latest performance.
Mask Autoencoder (MAE) MAE introduces two key hyperparameters: training strategies (linear detection vs. fine-tuning) and masking rate, which are explored over a range: $[0.10,0.90]$ with an interval of 0.1. The masking rate directly affects the difficulty of the reconstruction task and the quality of the learning representation, while the training strategy determines how to adapt to the downstream task. In our experiments (see Supplementary Table 9 for more details), GPT-4o-mini achieved explicit performance with efficiency, demonstrating the extraordinary ability to identify the best mask configuration and training strategies. Qwen2.5-7B achieves early convergence and disciplined stopping. The DeepSeek-R1 series shows consistent scaling behavior: performance improves from 7B to 14B to 32B, but slightly drops at 70B. In traditional solvers, GS and BS lag significantly behind LLM-based solvers, highlighting the advantages of intelligent hyperparameter optimization methods. Apart from Qwen2.5-14b and DeepSeek-R1-7b, most solvers can find the optimal solution, but significant differences exist among the various methods in terms of search efficiency and convergence speed.
ResNet Family On the ImageNet ResNet hyperparameter benchmark, LLMs in the 7B–32B range demonstrate superior optimization efficacy. The top-performing solvers, such as Qwen2.5-14B (1st), DeepSeek-R1-32B (2nd), and Qwen2.5-7B with DeepSeek-R1-14B (tied 3rd), consistently identify the optimal ResNet configuration within one or two trials, halting immediately with near-perfect sample efficiency. In contrast, larger variants from the same model families drop to middle or lower rankings, consuming over half the search budget with notably higher exploration overhead. While GPT-series models, as large-scale counterparts, still achieve competitive results, the 7B–32B class exhibits a clear advantage in balancing accuracy and efficiency. For the architectural search on ImageNet, we explored key dimensions such as depth and bottleneck design. Qwen2.5-14B attains theoretical peak performance, with its 7B and 32B versions completing further exploration efficiently. These results reflect an ability to identify ideal network configurations with minimal wasted exploration.
LCBench We evaluate SelfAI on the standard AutoML benchmark [71] using Bayesian optimization over 2000 rounds of hyperparameter search, covering critical parameters including learning rate, batch size, network depth, and dropout rate. As shown in Supplementary Table 11, DeepSeek-R1-70B, Qwen2.5-14B, and Llama3.3-70B secure the top three positions in the ranking, with DeepSeek-R1-70B attaining the highest overall score and the lowest search diversity. Notably, leading models in the 7B–32B parameter range efficiently identify near-optimal hyperparameter configurations by evaluating only a minimal set of candidate solutions, achieving an order-of-magnitude improvement in stop-time compared to classical methods. These findings demonstrate diminishing returns with scaling, as larger models within the same series (e.g., Qwen2.5-72B) typically underperform comparable models of medium size, while the 7B model can achieve robust performance comparable to models several times its size.

### v2

@@ 1 Introduction
Addressing these limitations requires a shift from execution-oriented autonomy toward cognitively oriented autonomy, in which discovery systems reason about exploration strategies, evaluate efficiency-diversity trade-offs, and adaptively decide when to stop. Motivated by this perspective, we introduce SelfAI, a self-directed, multi-agent-enabled discovery system that treats scientific discovery as a trajectory-level decision-making process. We integrate human intent, exploration reasoning, and experimental execution into a unified closed-loop framework. These are built with three modules and external tools (Fig. 1a), enabling efficiency-diversity exploration in long-horizon scientific workflows. To evaluate discovery quality, we introduce two novel complementary metrics, Score (Discovery Efficiency) and $\text{AUP}_{D}$ (Area Under the Performance Diversity). Score aggregates, across tasks, the normalized improvement over the search space together with penalties for discovering good configurations late and for stopping far from the best-found point. $\text{AUP}_{D}$ explicitly encodes how broadly a solver explores the search space by summarizing the performance-diversity tradeoffs across the entire trajectory, enabling detailed analysis of exploration behavior and stopping decisions in long-term searches. Together, these metrics measure quantitative assessments of exploration behavior, reasoning structure, and stopping decisions in long-term autonomous experiments.
Figure 1: SelfAI Framework for Automated Scientific Experimentation. a, Holistic architecture of the multi-agent system, which transforms various experiments in the research process into a structured workflow. b, User intents, comprising ideas and experiment schemes, are transformed into structured configurations via a predefined prompt. These inputs are processed through successive stages: hypothesis generation, strategic planning, trial execution, and result collection.
As illustrated in Fig. 1b, SelfAI operates as a closed-loop discovery system that unifies research intent, trajectory-aware reasoning, and a diverse set of executive experiment tools to support various scientific problems and long-horizon scientific discovery. The workflow begins with a User Agent, which translates high-level research objectives and exploratory questions into standardized experiment configurations. These structured configurations provide a consistent interface across heterogeneous environments, enabling reliable comparison and reuse of results. Cognitive Agent evaluates trial outcomes, reasons over accumulated discovery trajectories, and adaptively revises exploration strategies over time. Experimental execution is coordinated by Experiment Manager, which autonomously manages resource allocation, environment provisioning, checkpoint-based execution, and comprehensive experiment logging. Through continual feedback between reasoning and execution, the Cognitive Agent and Experiment Manager jointly assess the expected utility of continued exploration and apply adaptive stopping decisions to terminate unproductive search paths. In this way, SelfAI enables sustained improvement through long-horizon, trajectory-driven discovery.
In summary, SelfAI provides a self-directed framework that treats scientific discovery as a strategic, trajectory-driven process. By iteratively integrating intent specification, trajectory-aware reasoning, and adaptive stopping, SelfAI enables principled allocation of experimental resources while preserving exploration diversity under realistic computational and experimental constraints. Across a broad range of scientific discovery tasks, SelfAI consistently achieves efficient, diverse, and high-quality exploration in heterogeneous and large-scale search spaces. These results demonstrate that trajectory-aware reasoning and adaptive stopping allow discovery systems to reduce redundant experimentation, identify diminishing returns early, and redirect effort toward more informative regions of the search space. More broadly, this work highlights the role of LLM-guided decision-making in organizing scientific exploration under explicit efficiency-diversity trade-offs.
Figure 2: Performance comparison and efficiency-diversity trade-offs across solvers. a, Normalized performance of five solvers (GS, BS, LLM, LLM-ES, and SelfAI) evaluated by Score (higher is better), $\text{AUP}_{D}$, Best-time, and Stop-time (lower is better). Bars indicate solver-level averages, with values in parentheses denoting the best-achieved or optimal reference performance. b, Outcome distributions across solvers. Left: kernel density estimates illustrating score concentration and convergence tendencies. Right: violin plots with overlaid individual trials (sample size n shown), capturing performance variability. The dashed horizontal line denotes the evaluation threshold separating low- and high-performance regimes; annotated statistics (red/blue) highlight solver behavior above and below this threshold, where broader exploration facilitates rapid escape from low-performance regions and tighter distributions in high-performance regions enable localized refinement.
2 Results
2.1 SelfAI: Long-horizon scientific discovery
SelfAI is a trajectory-aware paradigm for long-horizon scientific discovery (Fig. 1). It treats accumulated experimental outcomes as structured evidence within an evolving discovery trajectory, from which exploration decisions, hypothesis refinement, and stopping judgments are iteratively and autonomously derived. SelfAI builds on recent LLM-based optimization and agentic frameworks that interleave reasoning with trial proposals but typically emphasize short-horizon improvement or final performance. Across experiments, SelfAI consistently synthesizes accumulated information, analyzes experimental trajectories, evaluates performance trends, and prioritizes promising regions of the search space.
2.2 Evaluation of long-horizon scientific discovery
To characterize efficiency-diversity trade-offs in long-horizon exploration, we propose four metrics: Score, $\text{AUP}_{D}$, $t_{\text{best}}$, and $t_{\text{stop}}$ (Fig. 2a), the details of which are described in Sect. 4.4. These metrics not only reflect final performance but also explicitly capture when high-quality solutions are discovered, how broadly the search space is explored, and how efficiently exploration is terminated. Score aggregates normalized improvement while penalizing both delayed discovery of optimal configurations and prolonged exploration after performance plateaus. $\text{AUP}_{D}$ (Area Under the Performance-Diversity Curve) quantifies the diversity of explored high-quality solutions and summarizes how performance gains are distributed along the discovery trajectory. The metrics $t_{\text{best}}$ and $t_{\text{stop}}$ denote the normalized time first to identify the best result and the time at which exploration terminates, respectively. An effective solver minimizes both quantities, achieving early discovery while avoiding unnecessary trials. In this framework, effective exploration strategies attain a high Score with a relatively low $\text{AUP}_{D}$, indicating efficient discovery with minimal wasted exploration. Detailed per-task comparisons of Score, $\text{AUP}_{D}$, $t_{\text{best}}$, and $t_{\text{stop}}$ are provided in the Supplementary Materials B.
2.3 SelfAI reorganizes exploration dynamics across scientific domains
We evaluate long-horizon scientific discovery using a benchmark constructed from real-world scientific discovery tasks (Materials and Methods) and assess solver performance under these efficiency-diversity trade-offs (Fig. 2). Using the same underlying language model (GPT-4o) across all solvers, we observe that while most methods are eventually able to identify high-performing configurations, their exploration dynamics differ substantially. As shown in Fig. 2a, SelfAI consistently outperforms LLM and LLM-ES baselines on $\text{AUP}_{D}$, Best-Time, and Stop-Time, indicating reduced redundant experimentation and more effective allocation of trials toward informative regions of the search space. Distributional analyses (Fig. 2b) further reveal that SelfAI reallocates experimental effort away from diminishing-return regions earlier while maintaining sufficient coverage of uncertain regions. Across the benchmark, for low-performance regimes, SelfAI obtains higher variances, conducts fewer trials to escape low-performance regimes, and concentrates effort in promising regions with lower variances. Finally, SelfAI achieves highly efficient exploration with 168 runs by terminating exploration near the best-achieved configuration under fixed trial budgets. Benchmark construction details and additional analyses are provided in Fig. 5.
2.3.1 Case studies
We present representative case studies to illustrate how trajectory-aware reasoning and adaptive stopping manifest in concrete scientific discovery scenarios, illustrating how different solvers organize exploration trajectories over time, allocate experimental effort, and determine when further experimentation becomes uninformative. In addition, we illustrate that trajectory-level metrics capture discovery efficiency, diversity, and how reasoning limitations can emerge in structured optimization settings. Additional task-specific analyses, including 12 tasks, are provided in Supplementary Sect. B.
Machine learning. We first evaluate SelfAI on the Boston housing price prediction task [boston] using a random forest regression model. As shown in Fig. 3 and Supplementary Table 4, nearly all solvers achieve a similar and high best result ($\approx 0.841$), indicating that differences in final predictive accuracy are minimal. Small- and medium-sized models outperform other solvers, but while large models DeepSeek-R1-70B and Llama3.3-70B demonstrated consistent and sustained exploration capabilities, their stopping behavior was less efficient, leading to longer search trajectories and delayed convergence. These results illustrate the importance of trajectory-aware reasoning and adaptive stopping in practical scientific optimization, as larger-scale models may exhibit either overly diffuse exploration or premature stopping and are not always aligned with task-specific objectives. Further examples are presented below in this section.
Computer vision. Fig. 4 illustrates these trajectory-level dynamics in the hyperparameter search landscape for image segmentation using SIREN (see also Supplementary Figs. 7-9). The smoothed response surface reveals several obvious peaks, indicating multiple competing high-performance regions. Starting from identical initial points, the tree-based Bayesian optimizer follows a spiral-like trajectory that underutilizes promising regions and exhibits limited global coverage, frequently converging to suboptimal local minima. In contrast, LLM-based solvers infer trends from accumulated evaluations and adapt exploration accordingly, resulting in broader coverage of the search space and more reliable identification of globally competitive configurations. Importantly, these solvers also monitor trajectory-level progress and terminate exploration once improvements plateau. This behavior directly translates into higher scores and lower $\text{AUP}_{D}$, demonstrating how trajectory-aware metrics capture both discovery efficiency and controlled diversity beyond final performance alone.
Medical image analysis. We introduce medical image segmentation using nnU-Net [nnunet], exploring performance optimization and best practices in medical image segmentation across 19 representative CNN-, Transformer-, and Mamba-based models. These results provide controlled settings for evaluating long-horizon optimization behavior. As shown in Fig. 3 and Supplementary Tables 12-13, trajectory-aware metrics reveal substantial differences in solver behavior that are not apparent from final segmentation accuracy alone. In particular, on the BraTS [brats] and BTCV [btcv] datasets, small- and medium-sized models achieve higher scores and lower $\text{AUP}_{D}$, reflecting their ability to rapidly identify high-performing configurations and terminate exploration. Trajectory-aware metrics concentrate exploration earlier in high-potential regions and avoid prolonged exploration once performance gains saturate. Conversely, solvers with comparable final performance but lower scores exhibit broader and more redundant exploration trajectories, leading to delayed stopping and reduced discovery efficiency.
Scientific computing. Tensor decomposition [hackbusch2014numerical, verstraete2008matrix] is a fundamental tool in scientific computing for high-dimensional data analysis, with broad applications in data compression [li2025lossless], computational acceleration [berezutskii2025tensor], and multimodal data fusion [steyaert2023multimodal]. The associated optimization problem [tw] presents a structured landscape characterized by competing objectives between compression quality and computational efficiency. Effective solvers must therefore balance refinement of promising configurations with controlled exploration of alternative regimes. As shown in Supplementary Table 6, GPT4-o3-mini performed well and ranked highly. However, Qwen2.5-72b may have over-relied on general reasoning ability and failed to balance mathematical knowledge and reasoning ability. Among the DeepSeek series models, although the DeepSeek-R1-7b model failed to find the optimal solution, its search strategy was acceptable, and its score was moderate. Although the DeepSeek-R1-70b and GPT-4o models have excellent search diversity and early stopping capabilities, their scores dropped. These results reveal that some large-scale models exhibit either overly diffuse exploration or premature stopping for task-specific objectives. This case illustrates that expressive reasoning capacity alone is insufficient for effective scientific discovery. Instead, successful optimization requires trajectory-level alignment between exploration strategy, stopping decisions, and problem structure. Supplementary analyses further document representative success and failure patterns across tasks (Supplementary Sect. B.6). In particular, hit rates for identifying globally optimal configurations vary substantially across models, with premature stopping limiting convergence in some cases. Moreover, increased model scale does not systematically improve long-horizon inference: larger models often sustain extended hypothesis exploration, leading to delayed commitment and greater performance variability under fixed experimental budgets. In contrast, several medium-sized models exhibit more stable trajectory organization, reinforcing the role of disciplined inference and adaptive stopping in efficient scientific discovery.
Result summary Across 12 heterogeneous tasks and language-model backbones (Supplementary Sect.  B), these results demonstrate that SelfAI yields consistent trajectory-level benefits. SelfAI does not merely improve convergence accuracy but fundamentally reorganizes exploration dynamics in long-horizon scientific discovery. By reallocating experimental effort away from diminishing-return regions toward structurally informative areas of the search space, SelfAI implements a policy-adaptive exploration strategy that aligns with identifiable principles of long-horizon scientific discovery.
2.4 Emergent principles of long-horizon scientific discovery
We further investigate whether long-horizon scientific discovery exhibits consistent, task-agnostic principles when guided by trajectory-aware reasoning. Fig. 4 illustrates these trajectory-level dynamics in the hyperparameter search landscape for image segmentation using SIREN (Sect. 2.3.1). From principles of long-horizon scientific discovery, trajectory-level dynamics follow a characteristic three-stage trajectory: rapid escape from low-efficiency regions, early concentration of effort in high-potential areas, and termination near the optimal configuration. This structured progression highlights the central role of strategic trial allocation and adaptive stopping in the efficiency of long-horizon exploration.
Aggregated performance across tasks and domains reveals additional trajectory-level regularities (Fig. 3). Despite task-dependent variations in relative rankings (Supplementary Fig. 6), SelfAI consistently achieves high Scores while maintaining relatively low $\text{AUP}_{D}$ values across heterogeneous tasks and language-model backbones. These results indicate that efficient discovery trajectories can be organized according to shared principles that are not tied to any single task or domain.
Model scale alone does not systematically improve long-horizon discovery behavior. Larger models often sustain prolonged exploration of alternative hypotheses, leading to delayed commitment and reduced adaptability under fixed experimental budgets. In contrast, small and mid-sized models often exhibit more stable cumulative progress, earlier concentration on informative regions, and more reliable stopping behavior. These observations show that baseline LLM solvers relying on exhaustive exploration without principled stopping often exhaust their search budgets (Fig. 1). In contrast, SelfAI-enabled solvers achieve substantially higher hit rates while terminating exploration significantly earlier. Notably, final performance is often near saturation across methods, with the best results differing only marginally. Nevertheless, significant differences exist in performance and diversity metrics, revealing substantial differences in identifying high-quality configurations and effectiveness in avoiding redundant experiments.
Figure 3: The top three rows present the Score of all solvers across different tasks, which assess the best stopping performance of SelfAI and baselines using different LLMs. The bottom three rows present the Diverse Metrics ($\text{AUP}_{D}$) of all solvers across different tasks, evaluating the diversity of solution trajectories.